OpenStudy (tamara7301):

i need help i have y=-2x+2 my teacher wants me to use the y2-y1 and x2-x1 but idk how plz help this is so confusing

OpenStudy (jango_in_dtown):

We have y=-2x+2 this gives y-2=-2x which gives y-2=(0-2)/(1-0)(x-0)

OpenStudy (tamara7301):

i dont understnd

OpenStudy (texaschic101):

ok....to use (y2 - y1) / (x2 - x1) you will need 2 sets of points. y = -2x + 2....we can already tell that the slope is -2....but if you want to use the slope formula, we can find you 2 points by finding the x and y intercepts. y = -2x + 2.....to find the y intercept, sub in 0 for x y = -2(0) + 2 y = 2....so one set of points is : (0,2) y = -2x + 2 ...to find the x intercept, sub in 0 for y 0 = -2x + 2 2x = 2 x = 1....so your x intercept is : (1,0) now lets sub it into the slope formula and see if the slope equals -2 as well. slope = (y2 - y1) / (x2 - x1) (0,2)....x1 = 0 and y1 = 2 (1,0)...x2 = 1 and y2 = 0 slope = (0 - 2) / (1 - 0) slope = -2/1 which is the same as -2 either way, you can find the slope...by using 2 sets of points and the slope formula...or turning your equation into y = mx + b form where m is your slope. questions ?

OpenStudy (tamara7301):

i dont think so

OpenStudy (texaschic101):

ok...well if you happen to think of any questions about this, just tag me and I will see what I can do :)

OpenStudy (tamara7301):

okay thx